OpenStudy (anonymous):

Solve for x. 6/x=x/12

OpenStudy (anonymous):

\[\frac{6}{x}=\frac{x}{12}\] Cross multiply! \[x^2=72\] \[x=\pm \sqrt{36*2}=\pm6\sqrt{2}\] Does that answer your question? :-)

OpenStudy (anonymous):

yes thanks